# Linear constellations in primes with arithmetic restrictions

### v2

@@ 6.3. Equidistribution in progressions
This does not follow directly from effective versions of the Chebotarev density theorem ([25, 38]) applied to $K(\mu_{q})$, as the degree $[K(\mu_{q}):\mathbb{Q}]$ can be too large in terms of $N$. Instead, we adapt techniques of Kane to our situation. We will use the following lemma, inspired by [23, Corollary 4 and Corollary 5], to identify the main term. We identify $(\mathbb{Z}/q\mathbb{Z})^{\times}\cong\operatorname{Gal}(\mathbb{Q}(\mu_{q})/\mathbb{Q})$ via the Artin symbol $\sigma_{a}=[\mathbb{Q}(\mu_{q})/\mathbb{Q},a]$ and may thus identify characters $\chi$ with characters $\chi^{\prime}$ of $\operatorname{Gal}(\mathbb{Q}(\mu_{q})/\mathbb{Q})$ via
@@
[…]
@@
[…]
@@
[…]
@@
As $[F/L,\mathfrak{p}]$ hits all conjugacy classes of $\operatorname{Gal}(F/L)$ by Chebotarev, we see that $\chi^{\prime}(\sigma)=\chi(\sigma|_{K})$ for all $\sigma\in\operatorname{Gal}(F/L)$, so $\chi^{\prime}$ is trivial on $\operatorname{Gal}(F/K)=\operatorname{Gal}(M/K\cap M)$. As $\chi^{\prime}$ is injective on $\operatorname{Gal}(M/\mathbb{Q})$, this shows that $K\cap M=M$, and thus $F=K$. Therefore, $\chi^{\prime}$ extends $\xi$ to $\operatorname{Gal}(K/\mathbb{Q})$. As $\chi^{\prime}$ is trivial on $\operatorname{Gal}(K/M)$, it is in particular trivial on $\operatorname{Gal}(K/K\cap\mathbb{Q}(\mu_{q}))$, which shows (2).
[…]
@@
If $q>(\log N)^{A+1}$, the statement is trivial due to (3.1). Hence, let us assume that $q\leq(\log N)^{A+1}$. We start from the formula in Lemma 6.2 with $F=1$. The inner sum over ideals $\mathfrak{a}$ of $\mathscr{O}_{L}$ on the right-hand side is just $F_{L,\chi\xi,N}(0)$ in the notation of [23, Definition 4]. Hence, [23, Proposition 14] yields the estimate
@@
For the second sum, we consider $\psi_{\xi}^{\prime}$ and $\xi$ as characters on $\operatorname{Gal}(K_{q}^{\operatorname{ab}}/\mathbb{Q})$, satisfying $\psi_{\xi}([K_{q}^{\operatorname{ab}}/\mathbb{Q},p])=\xi([K_{q}^{\operatorname{ab}}/\mathbb{Q},p])$ for all primes $p\nmid q\Delta_{K}$, so by Chebotarev they agree on all of $\operatorname{Gal}(K_{q}^{\operatorname{ab}}/\mathbb{Q})$. Observe that the image of the map in Lemma 6.4, (1), is equal to $\operatorname{Gal}(K_{q}^{\operatorname{ab}}/L\cap\mathbb{Q}(\mu_{q}))$. If $\sigma_{b}|_{\mathbb{Q}(\mu_{q})\cap L}=1$, then both $c|_{K_{q}^{\operatorname{ab}}}$ and $\sigma_{b}|_{K_{q}^{\operatorname{ab}}}$ are contained in this image, and by character orthogonality inner sum over $\xi$ becomes
@@
$\mathbb{E}_{n\in[N]}\mathbf{1}_{P}(n)\left(\Lambda_{K,C,b,W}(n)-\phi(W)\eta_{K,C}(b,W)\Lambda_{b,W}(n)\right)\ll_{K,A}(\log N)^{-A}.$
@@
$\mathbb{E}_{n\in[N]}\mathbf{1}_{P}(n)\Lambda_{K,C,b,W}(n)$ $=\frac{\phi(W)}{NW}\sum_{m=1}^{\tilde{N}}\Lambda_{K,C}((b+lW)+mqW)=\frac{\phi(W)}{NW}\sum_{\begin{subarray}{c}b+lW<n\leq b+lW+q\tilde{N}W\\ n\equiv b+lW\bmod qW\end{subarray}}\Lambda_{K,C}(n).$
@@
$\mathbb{E}_{n\in[N]}\mathbf{1}_{P}(n)\Lambda_{K,C,b,W}(n)$ $=\frac{\phi(W)q\tilde{N}}{N}\eta_{K,C}(b+lW,qW)+O_{K,A}\left((\log N)^{-A}\right).$ (6.5)
@@
$=\frac{\phi(W)q\tilde{N}}{N}\frac{\mathbf{1}_{(b,qW)=1}}{\phi(qW)}+O_{A}((\log N)^{-A}).$
@@
The proof that Lemma 6.6 implies Theorem 2.3 is almost the same as the proof, given on [14, pp.544–547], that [14, Proposition 2.1] implies [14, Theorem 1.1], with $\mu(n)$ replaced by our function $\Lambda_{K,C,b,W}(n)-\phi(W)\eta_{K,C}(b,W)\Lambda_{b,W}(n)$. The key differences are:
@@
    This is relevant in the deduction of [14, (2.6)], but the difference can be absorbed immediately by adapting the value of $A$.
  • Instead of [14, Proposition A.2] for $\mu(n)$, we use Lemma 6.5 to bound correlations as in [14, (2.9)] when $F_{j,k}$ is constant.
  • Instead of the application of [14, Proposition 2.1] at the very end of the proof, we apply Lemma 6.6.
@@ 7.1. The fields $F(q,k,a)$
([31, Lemma 2.3],[10, Lemma 2.2]) For $k$ square-free let $k^{\prime}=k/\gcd(k,h_{a})$. Then $[F(q,k,a):\mathbb{Q}]=k^{\prime}\phi(\operatorname{lcm}(q,k))/\epsilon(q,k)$, where
@@
([31, Lemma 2.4],[10, Lemma 2.4]) We have
$[\mathbb{Q}(\zeta_{q})\cap G(k,a):\mathbb{Q}(\zeta_{\gcd(q,k)})]=\begin{cases}2&\text{ if }2\mid k,\ \Delta_{a}\nmid k\text{ and }\Delta_{a}\mid\operatorname{lcm}(q,k),\\ 1&\text{ otherwise.}\end{cases}$
@@
$\mathbb{E}_{n\leq N}\Lambda_{a,b,W}^{\prime}(n)f(b+nW)=\sum_{\begin{subarray}{c}k\in\mathbb{N}\\ p_{+}(k)\leq w(N)\end{subarray}}\mu(k)\mathbb{E}_{n\leq N}\Lambda_{G(k,a),\{\operatorname{id}\},b,W}^{\prime}(n)f(b+nW)+O_{a,C,\mathscr{D}}\left(\frac{1}{w(N)}\right).$ (7.1)
@@
$\mathbb{F}_{p}^{\times}=\langle a\rangle\quad\Longleftrightarrow\quad p\nmid a\text{ and }\forall q\mid p-1,a\notin\mathbb{F}_{p}^{\times q}\quad\Longleftrightarrow\quad p\nmid a\text{ and }R_{a}(q,p)\text{ fails for all }q.$ (7.2)
For any $11\leq\eta_{1}\leq\eta_{2}\leq WN-$ 1, let
@@
Thus, the conditional effective Chebotarev theorem [34, Théorème 4] for $F(W_{0},q,a)$ yields the asymptotic
@@
From the $1$-boundedness of $F$, we moreover see that
@@ 9.1. Pseudorandom majorant for $\Lambda_{a,b,W}^{\prime}(n)$
Let $t\geq 1$ and $D>1$. Then there is a constant $C_{0}=C_{0}(D)$, such that the following holds. Let $C\geq C_{0}$ and let $N^{\prime}\in[CN,2CN]$ be prime. Let $a_{1},\ldots,a_{t}\in\mathbb{Z}$, such that no $a_{i}$ is equal to $-1$ or a perfect square. Let $b_{0},\ldots,b_{t}\in\{0,\ldots,W-1\}$ be coprime to $W$. Then there exists a function $\nu:\mathbb{Z}/N^{\prime}\mathbb{Z}\to(0,\infty)$ that has the following properties:
@@
For all $\tilde{s},\tilde{t}\in[D]$ and finite complexity systems $\tilde{\Psi}=(\tilde{\psi}_{1},\ldots,\tilde{\psi}_{\tilde{t}})$ of affine-linear forms on $\mathbb{Z}^{\tilde{s}}$ with the coefficients of all linear terms bounded in absolute value by $D$, we have
@@
A function $\nu$ satisfying this is constructed, essentially, in the proof of [13, Proposition 6.4] in [13, Appendix D]. The only difference to our situation is a minor change in the definition of $W$. In [13], the function $w(N)=\log\log\log N$ is fixed and $W=\prod_{p\leq w(N)}p$, whereas for us $w(N)$ can be any function that grows to infinity with $N$ sufficiently slowly, and $W$ is defined as in (2.1).
The proof in [13, Appendix D], with the obvious necessary modifications, goes through in our situation as well. See also [4] for a similar situation. Note that this leads to the dependence of the error term in the linear forms condition on $\mathscr{D}$ and $w(\cdot)$. Note, moreover, that $C_{0}=C_{0}(D)$ is independent of $\mathscr{D},W$, and the same applies to the bound in (9.1).
The majorant constructed in [13, Appendix D] also satisfies a so-called correlation condition. This condition was used in [13] for transferring the inverse theorem for Gowers norms, but is no longer required due to recent advances. ∎
@@
The case where $w(N)=\log\log\log N$ and $\mathscr{D}=1$ was proved by Green and Tao in [13, Theorem 7.2], and their proof also works in our situation. As a thorough verification of this claim places quite a burden on the reader, we will now also derive the lemma rigorously from Bienvenu’s result [4, Theorem 1.3], concerning affine-linear correlations of the von Mangoldt function in which the linear coefficients are allowed to grow logarithmically in $N$. Let us assume $\gcd(b,W)=1$, as otherwise the result is obvious. Then
@@ 9.3. Application of the generalised von Neumann theorem
The following lemma is essentially [13, Proposition 7.1]. By $\kappa_{a}(\delta)$, we denote a quantity that goes to zero with $\delta$, i.e. a quantity that is smaller in absolute value than any $\epsilon>0$ if only $\delta$ is sufficiently small in terms of $\epsilon$ and $a$.
@@
for all $\tilde{s},\tilde{t}\in[D]$ and all finite complexity systems $\tilde{\Psi}=(\tilde{\psi}_{1},\ldots,\tilde{\psi}_{\tilde{t}})$ of affine-linear forms on $\mathbb{Z}^{\tilde{s}}$ in which the coefficients of all linear terms are bounded in absolute value by $D$, and where $z:\mathbb{N}\to(0,\infty)$ is a function satisfying $\lim_{N\to\infty}z(N)=0$.
@@
The statement is almost the same as in [13, Proposition 7.1], except for two differences. Firstly, we do not suppose that the majorant $\nu$ satisfy a $D$-correlation condition. This is irrelevant, as the proof of [13, Proposition 7.1] given in [13, Appendix C] uses only the $D$-linear forms condition. Secondly, we state more explicitly the dependence of the estimate on the error term $z(N)$ in the linear forms condition, in particular that the term $\kappa_{C,s,t,u,L}(\delta)$ in the final estimate does not depend on this error term. One can see this by following carefully the proof in [13, Appendix C]. ∎

### v3

@@ 6.3. Equidistribution in progressions
This does not follow directly from effective versions of the Chebotarev density theorem ([27, 42]) applied to $K(\mu_{q})$, as the degree $K(\mu_{q}):\mathbb{Q}]$ can be too large in terms of $N$. Instead, we adapt techniques of Kane to our situation. We will use the following lemma, inspired by [24, Corollary 4 and Corollary 5], to identify the main term. We identify $(\mathbb{Z}/q\mathbb{Z})^{\times}\cong\operatorname{Gal}(\mathbb{Q}(\mu_{q})/\mathbb{Q})$ via the Artin symbol $\sigma_{a}=[\mathbb{Q}(\mu_{q})/\mathbb{Q},a]$ and may thus identify characters $\chi$ with characters $\chi^{\prime}$ of $\operatorname{Gal}(\mathbb{Q}(\mu_{q})/\mathbb{Q})$ via
@@
We write $K_{q}^{\operatorname{ab}}:=K\cap\mathbb{Q}(\mu_{q})$ and start with $\emph{(1)}\Rightarrow\emph{(2)}$. If (1) holds, we can consider $\xi$ as a character on the image of $\operatorname{Gal}(K/L)$ in $\operatorname{Gal}(K_{q}^{\operatorname{ab}}/\mathbb{Q})$, which is abelian, and thus extend it to a character on $\operatorname{Gal}(K_{q}^{\operatorname{ab}}/\mathbb{Q})$. Via restriction $\operatorname{Gal}(K/\mathbb{Q})\to\operatorname{Gal}(K_{q}^{\operatorname{ab}}/\mathbb{Q})$, we may view this extension as a character on $\operatorname{Gal}(K/\mathbb{Q})$ trivial on $\operatorname{Gal}(K/K_{q}^{\operatorname{ab}})$, which shows (2).
The implication $\emph{(2)}\Rightarrow\emph{(1)}$ is trivial. To prove $\emph{(2)}\Rightarrow\emph{(3)}$, we consider $\xi$ as a character of $\operatorname{Gal}(K_{q}^{\operatorname{ab}}/\mathbb{Q})$, so by restriction it induces a character $\psi^{\prime}$ of $\operatorname{Gal}(\mathbb{Q}(\mu_{q})/\mathbb{Q})$ that is trivial on $\operatorname{Gal}(\mathbb{Q}(\mu_{q})/K_{q}^{\operatorname{ab}})$. For the corresponding Dirichlet character $\psi$ and every prime $p\nmid q\Delta_{K}$, we then have
@@
For $\emph{(3)}\Rightarrow\emph{(4)}$, we show that $\psi$ is such a Dirichlet character. Let $\mathfrak{p}$ be a prime ideal of $\mathscr{O}_{L}$ not dividing $q\Delta_{K}$, and write $\mathfrak{N}\mathfrak{p}=p^{f}$ for some prime $p\nmid q\Delta_{K}$ and $f\in\mathbb{N}$. By the formalism for the Artin symbol,
@@
For $\emph{(4)}\Rightarrow\emph{(2)}$, let $\chi$ be a Dirichlet character as in (3). Let $M\subseteq\mathbb{Q}(\mu_{q})$ be the fixed field of the kernel of $\chi^{\prime}$, so we can consider $\chi^{\prime}$ as an injective character of $\operatorname{Gal}(M/\mathbb{Q})$. Consider the compositum $F=KM$. Via restriction $\operatorname{Gal}(F/\mathbb{Q})\to\operatorname{Gal}(M/\mathbb{Q})$, we may now consider $\chi^{\prime}$ as a character of $\operatorname{Gal}(F/\mathbb{Q})$ that is trivial on $\operatorname{Gal}(F/M)$. For any prime ideal $\mathfrak{p}$ of $L$ not dividing $q\Delta_{K}$ with $\mathfrak{N}\mathfrak{p}=p^{f}$, we have
@@
As $F/L,\mathfrak{p}]$ hits all conjugacy classes of $\operatorname{Gal}(F/L)$ by Chebotarev, we see that $\chi^{\prime}(\sigma)=\chi(\sigma|_{K})$ for all $\sigma\in\operatorname{Gal}(F/L)$, so $\chi^{\prime}$ is trivial on $\operatorname{Gal}(F/K)=\operatorname{Gal}(M/K\cap M)$. As $\chi^{\prime}$ is injective on $\operatorname{Gal}(M/\mathbb{Q})$, this shows that $K\cap M=M$, and thus $F=K$. Therefore, $\chi^{\prime}$ extends $\xi$ to $\operatorname{Gal}(K/\mathbb{Q})$. As $\chi^{\prime}$ is trivial on $\operatorname{Gal}(K/M)$, it is in particular trivial on $\operatorname{Gal}(K/K\cap\mathbb{Q}(\mu_{q}))$, which shows (2).
Now suppose that (1)(4) hold. For the additional assertion, let $\rho$ be a Dirichlet character modulo $q$. We need to show that $\rho(\mathfrak{N}\mathfrak{p})=1$ for all $\mathfrak{p}\nmid q\Delta_{K}$ if and only if $\rho^{\prime}$ is trivial on $\operatorname{Gal}(\mathbb{Q}(\mu_{q})/L\cap\mathbb{Q}(\mu_{q}))$. This is similar to the proof of $\emph{(4)}\Rightarrow\emph{(2)}$ above and is shown in [24, Corollary 5]. ∎
@@
If $q>(\log N)^{A+1}$, the statement is trivial due to (3.1). Hence, let us assume that $q\leq(\log N)^{A+1}$. We start from the formula in Lemma 6.2 with $F=1$. The inner sum over ideals $\mathfrak{a}$ of $\mathscr{O}_{L}$ on the right-hand side is just $F_{L,\chi\xi,N}(0)$ in the notation of [24, Definition 4]. Hence, [24, Proposition 14] yields the estimate
@@
For the second sum, we consider $\psi_{\xi}^{\prime}$ and $\xi$ as characters on $\operatorname{Gal}(K_{q}^{\operatorname{ab}}/\mathbb{Q})$, satisfying $\psi_{\xi}([K_{q}^{\operatorname{ab}}/\mathbb{Q},p])=\xi([K_{q}^{\operatorname{ab}}/\mathbb{Q},p])$ for all primes $p\nmid q\Delta_{K}$, so by Chebotarev they agree on all of $\operatorname{Gal}(K_{q}^{\operatorname{ab}}/\mathbb{Q})$. Observe that the image of the map in Lemma 6.4, (1), is equal to $\operatorname{Gal}(K_{q}^{\operatorname{ab}}/L\cap\mathbb{Q}(\mu_{q}))$. If $\sigma_{b}|_{\mathbb{Q}(\mu_{q})\cap L}=1$, then both $c|_{K_{q}^{\operatorname{ab}}}$ and $\sigma_{b}|_{K_{q}^{\operatorname{ab}}}$ are contained in this image, and by character orthogonality, the inner sum over $\xi$ becomes
@@
$\mathbb{E}_{n\in[N]}\mathbf{1}_{P}(n)\left(\Lambda_{K,C,b,W}(n)-\phi(W)\eta_{K,C}(b,W)\Lambda_{b,W}(n)\right)\ll_{K,A,\mathscr{D}}(\log N)^{-A}.$
@@
$\mathbb{E}_{n\in[N]}\mathbf{1}_{P}(n)\Lambda_{K,C,b,W}(n)$ $=\frac{\phi(W)}{NW}\sum_{m=1}^{\tilde{N}}\Lambda_{K,C}((b+lW)+mqW)=\frac{\phi(W)}{NW}\hskip-14.22636pt\sum_{\begin{subarray}{c}b+lW<n\leq b+lW+q\tilde{N}W\\ n\equiv b+lW\bmod qW\end{subarray}}\hskip-8.5359pt\Lambda_{K,C}(n).$
@@
$\mathbb{E}_{n\in[N]}\mathbf{1}_{P}(n)\Lambda_{K,C,b,W}(n)$ $=\frac{\phi(W)q\tilde{N}}{N}\eta_{K,C}(b+lW,qW)+O_{K,A,\mathscr{D}}\left((\log N)^{-A}\right).$ (6.5)
@@
$=\frac{\phi(W)q\tilde{N}}{N}\frac{\mathbf{1}_{(b+lW,qW)=1}}{\phi(qW)}+O_{A,\mathscr{D}}((\log N)^{-A}).$
@@
The proof that Lemma 6.6 implies Theorem 2.3 is almost the same as the proof, given on [15, pp.544–547], that [15, Proposition 2.1] implies [15, Theorem 1.1], with $\mu(n)$ replaced by our function $\Lambda_{K,C,b,W}(n)-\phi(W)\eta_{K,C}(b,W)\Lambda_{b,W}(n)$. The key differences are:
@@
    This is relevant in the deduction of [15, (2.6)], but the difference can be absorbed immediately by adapting the value of $A$.
  • Instead of [15, Proposition A.2] for $\mu(n)$, we use Lemma 6.5 to bound correlations as in [15, (2.9)] when $F_{j,k}$ is constant.
  • Instead of the application of [15, Proposition 2.1] at the very end of the proof, we apply Lemma 6.6.
@@ 7.1. The fields $F(q,k,a)$
([35, Lemma 2.3],[10, Lemma 2.2]) For $k$ square-free let $k^{\prime}=k/\gcd(k,h_{a})$. Then $F(q,k,a):\mathbb{Q}]=k^{\prime}\phi(\operatorname{lcm}(q,k))/\epsilon(q,k)$, where
@@
([35, Lemma 2.4],[10, Lemma 2.4]) We have
[…]
@@
$\mathbb{E}_{n\leq N}\Lambda_{a,b,W}^{\prime}(n)f(b+nW)=\hskip-8.5359pt\sum_{\begin{subarray}{c}k\in\mathbb{N}\\ p_{+}(k)\leq w(N)\end{subarray}}\hskip-8.5359pt\mu(k)\mathbb{E}_{n\leq N}\Lambda_{G(k,a),\{\operatorname{id}\},b,W}^{\prime}(n)f(b+nW)+O_{a,C,\mathscr{D}}\left(\frac{1}{w(N)}\right).$ (7.1)
@@
[…]
For any $1\leq\eta_{1}\leq\eta_{2}\leq WN-1$, let
@@
Thus, the conditional effective Chebotarev theorem [38, Théorème 4] for $F(W_{0},q,a)$ yields the asymptotic
@@
From the 11-boundedness of $F$, we moreover see that
@@ 9.1. Pseudorandom majorant for $\Lambda_{a,b,W}^{\prime}(n)$
Let $t\geq 1$ and $D>1$. Then there is a constant $C_{0}=C_{0}(D)$, such that the following holds. Let $C\geq C_{0}$ and let $N^{\prime}\in[CN,2CN]$ be prime. Let $a_{1},\ldots,a_{t}\in\mathbb{Z}$, such that no $a_{i}$ is equal to 1-1 or a perfect square. Let $b_{0},\ldots,b_{t}\in\{0,\ldots,W-1\}$ be coprime to $W$. Then there exists a function $\nu:\mathbb{Z}/N^{\prime}\mathbb{Z}\to(0,\infty)$ that has the following properties:
@@
[…]
@@
A function $\nu$ satisfying this is constructed, essentially, in the proof of [14, Proposition 6.4] in [14, Appendix D]. The only difference to our situation is a minor change in the definition of $W$. In [14], the function $w(N)=\log\log\log N$ is fixed and $W=\prod_{p\leq w(N)}p$, whereas for us $w(N)$ can be any function that grows to infinity with $N$ sufficiently slowly, and $W$ is defined as in (2.1).
The proof in [14, Appendix D], with the obvious necessary modifications, goes through in our situation as well. See also [4] for a similar situation. Note that this leads to the dependence of the error term in the linear forms condition on $\mathscr{D}$ and $w(\cdot)$. Note, moreover, that $C_{0}=C_{0}(D)$ is independent of $\mathscr{D},W$, and the same applies to the bound in (9.1).
The majorant constructed in [14, Appendix D] also satisfies a so-called correlation condition. This condition was used in [14] for transferring the inverse theorem for Gowers norms, but is no longer required due to recent advances. ∎
@@
The case where $w(N)=\log\log\log N$ and $\mathscr{D}=1$ was proved by Green and Tao in [14, Theorem 7.2], and their proof also works in our situation. As a thorough verification of this claim places quite a burden on the reader, we will now also derive the lemma rigorously from Bienvenu’s result [4, Theorem 1.3], concerning affine-linear correlations of the von Mangoldt function in which the linear coefficients are allowed to grow logarithmically in $N$. Let us assume $\gcd(b,W)=1$, as otherwise the result is obvious. Then
@@ 9.3. Application of the generalised von Neumann theorem
The following lemma is essentially [14, Proposition 7.1]. By $\kappa_{a}(\delta)$, we denote a quantity that goes to zero with $\delta$, i.e. a quantity that is smaller in absolute value than any $\epsilon>0$ if only $\delta$ is sufficiently small in terms of $\epsilon$ and $a$.
@@
[…]
@@
The statement is almost the same as in [14, Proposition 7.1], except for two differences. Firstly, we do not suppose that the majorant $\nu$ satisfy a $D$-correlation condition. This is irrelevant, as the proof of [14, Proposition 7.1] given in [14, Appendix C] uses only the $D$-linear forms condition. Secondly, we state more explicitly the dependence of the estimate on the error term $z(N)$ in the linear forms condition, in particular that the term $\kappa_{C,s,t,u,L}(\delta)$ in the final estimate does not depend on this error term. One can see this by following carefully the proof in [14, Appendix C]. ∎